One less than certain positive number multiplied by three times this number less two is 14. Determine the number
Answer by richwmiller(17219)

One less than certain positive number
x-1
multiplied by three times this number *3x
less two -2
is 14. =14

Determine the number.
(x-1)*(3x-2)=14
3 x^2-5 x+2 = 14
x=-4/3 x=3
(3-1)(3*3-2)=14
2*7=14
-4/3 is negative
or
(x-1)*(3x)-2=14
has no integer solutions